OpenStudy (anonymous):

8. Essay: Show all work. A gardener wants to create a rectangular garden with length of 3x - 3y ft. and width of x - 5y ft. What is an algebraic expression for the area of the garden? Be sure to multiply this out, express in simplest correct mathematical form, and include units. (Points : 6)

OpenStudy (anonymous):

Length = 3x -3y ft Width = x -5y ft Area = length * width = (3x -3y)*(x-5y) sq. ft 3x*(x -5y) -3y*(x-5y) sq. ft 3x^2 -15xy -3yx +15y^2 3x^2 -18xy+15y^2 3(x^2-6xy+5y^2) sq. ft